OpenStudy (anonymous):

I'm having problem confirming they are inverses can someone please help me and explain each step?

OpenStudy (anonymous):

\[f(x)=\frac{ x-7 }{ x+3 } \]

OpenStudy (anonymous):

\[g(x)=\frac{ -3x-7 }{ x-1 }\]

OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x.

OpenStudy (anonymous):

Okay. Have you computed f(g(x)) yet?

OpenStudy (anonymous):

Yes and i got stuck at the bottom

OpenStudy (anonymous):

i got up to here:

OpenStudy (anonymous):

\[\frac{ -3x ^{2}-11x+14 }{ -3x ^{2}-x+4 }\]

OpenStudy (anonymous):

I would recheck your computation.

OpenStudy (amistre64):

what is g(0) and f(0)?

OpenStudy (anonymous):

@amistre64 im not sure what you're asking @RBauer4 okay give me a sec

OpenStudy (anonymous):

@RBauer4 i don't know what i did wrong

OpenStudy (amistre64):

was just wondering something, let x=0 prolly confusing reciprocal for inverse an easy check is to take say f(x) and swap out x for y, and solve for y to see if we get to g(x)

OpenStudy (anonymous):

Can you set it up please. I don't understand what you mean

OpenStudy (amistre64):

\[y=f(x)=\frac{ x-7 }{ x+3 }\] \[x=f(y)=\frac{y-7 }{y+3 }\] see if f(y) = g(x) by solving for y \[x=\frac{y-7 }{y+3 }\] \[x(y+3)=y-7\] \[yx+3x=y-7\] \[yx-y=-3x-7\] \[y(x-1)=-3x-7\] \[y=\frac{-3x-7}{x-1}\]

OpenStudy (anonymous):

@amistre64 I understand now, but I have to plug it in the other way because I have to show the thought process

OpenStudy (anonymous):

Note, the question does state to show f(g(x)) = x and that g(f(x)) = x.

OpenStudy (amistre64):

yeah :) but this was just an idea .... the rest is just algebra methods

OpenStudy (amistre64):

actually, this method shows that f(f(y)) = f(g(x)) = x in a longer format prolly

OpenStudy (anonymous):

Certainly, and I feel that the method of computing an inverse function is more meaningful than showing f(g(x)) = x and g(f(x)) = x; but, that's just my opinion.

OpenStudy (anonymous):

Can either of you help me doing it the other way... PLEASE

OpenStudy (anonymous):

\[f(g(x)) = \frac{ \frac{ -3x-7 }{ x - 1} - 7}{ \frac{ -3x - 7 }{ x-1 } + 3 }\]

OpenStudy (anonymous):

Multiply -7 by (x-1)/(x-1) and 3 by (x-1)/(x-1)

OpenStudy (amistre64):

or, mutliply top and bottom by (x-1) to clear out the fractiony looking messes :)

OpenStudy (anonymous):

i